OpenStudy (anonymous):

my teacher said 'if sec(x)=sec(y), then x=y' is a false statement why is this?

OpenStudy (anonymous):

because sec is a periodic function. so it can take on many values of x and y which results in the same sec value.

OpenStudy (amistre64):

..... yep, one value can have many names

OpenStudy (anonymous):

so we can say x=y+2npi for n is an integer right? would that make the statement true?

OpenStudy (anonymous):

The secant function is not injective so there is no unique inverse function. ;)

OpenStudy (anonymous):

sec(x) = sec(y) is the same as cos(x) = cos(y) and cos is has a period of 2pi. so yes.

OpenStudy (amistre64):

if we put john on a merry-go-round and turn him pi/2 ; and take billy and turn him 5pi/2 the sec(john) = sec(billy) ; even tho john does not equal billy

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

cot(x)*tan(x)=1 is false since cot(x)=cos(x)/sin(x) and sin(x) is sometimes 0 just like tan(x)=sin(x)/cos(x) and cos(x) is sometimes 0

OpenStudy (anonymous):

seriously monkey131? I don't quite agree with you on that hmm.

OpenStudy (anonymous):

the problem is that when sin(x) is 0 and blahblah there will always be a denominator with a 0. those values are undefined for the tangent/cotangent function.

OpenStudy (anonymous):

well we can't have divsion by 0 so it would be true if they said whenever cot(x) and tan(x) is not 0

OpenStudy (anonymous):

when sin(x) = 0 the value of tan(x) is not defined. so it wouldn't be considered as a falsified statement.

OpenStudy (anonymous):

i mean cos(x).

OpenStudy (anonymous):

both when both are 0 the equality fails

OpenStudy (anonymous):

you can't have both being 0 at the same time.

OpenStudy (anonymous):

right and you can't have 0=sin(x) and 0=cos(x) but also if sin(x)=0 or cos(x)=0 the equality fails but i guess and wouldn't happen sin(x) is not zero the sametime cos(x) is 0

OpenStudy (anonymous):

so they would never be zero at the same time so if just one of them is zero then the equality fails

OpenStudy (anonymous):

yup that's true. but you're assuming that 0 * infinity is not 1. what is 1/0 * 0 for all you know it could be 1!

OpenStudy (anonymous):

ok you lost me there

OpenStudy (anonymous):

ok you lost me there

OpenStudy (anonymous):

if tan(x) = 0 cot(x) = 1/0

OpenStudy (anonymous):

right i was say we can't have cos(x)=0 and we can't have sin(x)=0

OpenStudy (anonymous):

we cannot have either case

OpenStudy (anonymous):

yeah that was what i said earlier. If that is not possible then the case where it is 0 is not possible and the statement isn't false.

OpenStudy (anonymous):

i mean statement isn't true.

OpenStudy (anonymous):

yeah that is what i was saying in the beginning

OpenStudy (anonymous):

O.o but you claimed that cot(x)*tan(x) is false

OpenStudy (anonymous):

cot(x)*tan(x) = 1 is false*

OpenStudy (anonymous):

yes it is false

OpenStudy (anonymous):

it is true when sin(x) does not equal 0 or when cos(x) does not equal 0

OpenStudy (anonymous):

but we cannot suppose that what is not given

OpenStudy (anonymous):

they cannot equal 0 since for those values where they are either tan(x)or cot(x) does not have a definite value.

OpenStudy (anonymous):

so the statement is false

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

it is true as far as all mathematical definitions is concerned.

OpenStudy (anonymous):

that cot(x)tan(x)=1?

OpenStudy (anonymous):

probably. define tan(90 degrees)

OpenStudy (anonymous):

cot(x)tan(x) does not always equal 1 as we pointed out above sometimes it is undefined for certain values like x=0 or x=pi/2 remember cot(x)=cos(x)/sin(x) so cot(pi/2) is undefined

OpenStudy (anonymous):

oops tan(pi/2) is undefind

OpenStudy (anonymous):

what is (1/0)*0 why can't it be 1?

OpenStudy (anonymous):

because its not 1

OpenStudy (anonymous):

thats an underiminate form a form that cannot be determined

OpenStudy (anonymous):

thats an underiminate form a form that cannot be determined